# arXiv:2603.02804v1[quant-ph] 03 Mar 2026

### 03 Mar 2026 v1

@@ 4.3.2 Numerical experiment for a practical scale QML benchmark
Finally, assuming the training of a practical QML model, we measured the mean training time per epoch and peak memory usage for $1,000$ samples of randomly initialized states. For the numerical settings, we set $b=10$ for Triton fused and $b=20$ for Triton fused (mem save), based on the previous numerical results with a batch size of one. In addition, we set the mini-batch size to $60$ for Triton fused and $85$ for Triton fused (mem save), to fully use the GPU memory capacity, reaching a peak memory usage of approximately $95$ GB, as shown in Table 1.
As a result, the execution time was $1,273$ sec for Triton fused and $1,111$ sec for Triton fused (mem save), indicating that Triton fused (mem save) was approximately $10\%$ faster than Triton fused. Thus, the strategy of storing state vectors in reduced precision accommodates an approximately $1.4$ times larger mini-batch size, thereby enabling more effective utilization of GPU parallel computation resources. This resulted in a throughput improvement, outweighing the overhead of type conversion.
Based on these results, for a $20$-qubit HEA model including $60,000$ parameters, we estimate the training time to be approximately $15$ to $20$ hours per epoch for a data set containing tens of thousands of training samples, such as MNIST or CIFAR-10. Moreover, since our implementation is compatible with PyTorch’s standard ecosystem, it is easy to use multi-GPU distributed training (e.g., via DistributedDataParallel in PyTorch), enabling even faster training. Therefore, our proposed method enables training deeper and larger-scale QML models within a realistic timeframe.
@@
Figure 9: (a) Peak memory usage and (b) Execution time, when varying $b$ for a $20$-qubit, $1000$-layer ($d=1,000$) HEA with a batch size of one
Table 1: Performance comparison for minimizing the sum of expectation values at the final state of an HEA with 2020-qubit, $1,000$ layers ($60,000$ parameters) for $1,000$ samples
@@
Triton Fused 60 10 1,273 94.456
Triton Fused(mem save) 85 20 1,111 95.378
@@ 5 Conclusion
Furthermore, by combining our proposed method with gradient checkpointing, we verified the practical training of a deep QML model. First, to optimize the checkpoint block size (i.e., the number of HEA layers to be included in the checkpoint block), we measured the mean execution time and memory usage with a batch size of one. The results showed that the execution time is almost constant regardless of the checkpoint block size, whereas the memory usage is highly consistent with our theoretical analysis. In addition, we found that memory usage can be reduced by $30\%$ by temporarily storing the state vectors as bfloat16 converted from float32 (complex64). Based on these results, we estimate that we could train a $20$-qubit QML model with $60,000$ parameters on tens of thousands of training samples, such as MNIST or CIFAR-10, within $15$ to $20$ hours per epoch. Also, since our implementation is compatible with standard PyTorch functions, we can further accelerate training by distributing samples over multiple GPUs using utilities in PyTorch, such as DistributedDataParallel. Therefore, our combined approach enables the training of deep QML models on large datasets within a realistic timeframe, which would previously require a supercomputer or a multi-GPU cluster.
@@
The code used in this study will be available at GitHub (https://github.com/puyokw/triton_qml).
@@ B.1 The implementation of fused CZ gates and fused CNOT gates
Since a CNOT gate or a CZ gate is a unitary matrix ($U=U^{\dagger}$) with no variational parameters, we can update the quantum states and adjoint states using the same kernel function for both the forward and backward paths. Here, we describe the implementation for the forward path.
Let $c$ be the index of the control qubit and $t$ be the index of the target qubit. As we explained in Sec. 3.1, applying a CZ gate acting on the $c$-th control qubit and the $t$-th target qubit corresponds to multiplying the probability amplitude $a_{k}$ by 1-1 only when both the $c$-th and $t$-th bits of the index $k$ are $1$. Similarly, for the implementation of fused multiple CZ gates, let $c_{j}$ and $t_{j}$ denote the control and target qubits for the $j$-th gate in the fused operation. Each thread first counts the number of CZ gates where both the $c_{j}$-th and $t_{j}$-th bits of the index $k$ are $1$. If this total count is odd, the thread multiplies the corresponding probability amplitude $a_{k}$ by 1-1.
Specifically, we implemented this using Triton [triton2019], as shown in Listing 1. First, we prepare a bitmask (cz_mask) where only the $c_{j}$-th and $t_{j}$-th bits are set to $1$. For each thread index (offset), we check if the bitwise-AND between the index and the mask (cz_mask & offset) equals the mask itself (cz_mask). By accumulating these boolean evaluation results across all individual CZ gates within the fused operation via an XOR operation to compute the parity, we can determine whether to flip the sign of the amplitude. By applying this calculation to all $2^{n}$ amplitudes, we can update the state vector for the fused CZ gates.
@@
Although we used fused CZ gates in this paper, the efficient implementation of fused CNOT gates is also possible. Here, we describe how to implement a fused CNOT gate with out-of-place operations, since, as we mentioned before, PyTorch’s automatic differentiation requires inputs to be invariant. Also, recall that applying CNOT gate corresponds to swapping the pair of amplitudes between $a_{\ldots 1_{c}\ldots 0_{t}\ldots}$ and $a_{\ldots 1_{c}\ldots 1_{t}\ldots}$ (or between $a_{\ldots 0_{t}\ldots 1_{c}\ldots}$ and $a_{\ldots 1_{t}\ldots 1_{c}\ldots}$), so we can implement it by tracking the index. We start from the original index assigned to each thread, and trace the index to obtain the destination index. Specifically, as described in Listing 2, for the $j$-th CNOT gate in the fused operation, if the $c_{j}$-th bit of the current index is $1$, we update the current index (out_idx) by applying a bitwise XOR operation between the current index and a mask (target_mask) that has only the $t_{j}$-th bit set to $1$. By repeating this procedure across all individual CNOT gates within the fused operation, we obtain the final destination index. By allocating a new vector and writing the original probability amplitude to the resulting destination index (out_idx), we can update the quantum states as an out-of-place operation.
@@ Appendix C Memory behavior analysis of the PyTorch native implementation using $100$-layer HEA
In this section, to complement the experiments in Sec. 4.3, we describe the results of a classical simulation for $100$ HEA layers to verify the memory behavior of the PyTorch native implementation, which could not be evaluated in the large-scale numerical experiments in Sec. 4.3 due to excessive memory requirements. First, we verify our analytical model described in Sec. 3.4, especially for the PyTorch native implementation, using a smaller model consisting of $100$ HEA layers with $20$ qubits and $6,000$ parameters. In this numerical experiment, we also performed the task of minimizing the sum of expectation values for the observable $O=IXYZ\ldots IXYZ$ (with $IXYZ$ repeated $5$ times) at the final states, made by applying the quantum circuit to random initial quantum states for a $20$-qubit system.
@@
since we obtain $\lceil\frac{l_{\text{var}}}{\alpha}\rceil=7$ by substituting $l_{\text{var}}=60,\alpha=9,d=100,M_{\text{sv}}=8\times 2^{20}$ into Eq. (7). From this theoretical peak memory usage, the optimal block size is $b^{*}=3.7$, and the measured minimum peak memory usage is $528$ MB at $b=4$. Also, the theoretical peak memory usage for Triton fused (mem save) is
@@
The optimal block size for Triton fused (mem save) is $b^{*}=5.3$, and the measured minimum peak memory usage is $406$ MB at $b=5$. Thus, we can see that the theoretical values for peak memory usage are highly consistent with the measured values for the Triton fused and Triton fused (mem save) implementations. Note that a slight discrepancy between theoretical and measured values is observed when memory usage is below 1 GB, primarily due to the memory overheads such as CUDA contexts.
@@
Comparing the training time for $1$ epoch: Here, we compare the training times among the three implementations. For the checkpoint setting, based on the results of the above numerical experiments, we set the checkpoint block size $b$ to $b=1$ (every HEA layer) for the PyTorch native implementation, $b=4$ for Triton fused, and $b=5$ HEA layers for Triton fused (mem save). We performed the task of minimizing the sum of expectation values for the observable $O=(IXYZ)^{\otimes 5}$ at the final state by applying $100$ HEA layers to $1,000$ randomly initialized quantum states with a $20$-qubit system over $1$ epoch.
As shown in Table 2, we set the mini-batch size to $60$ for the PyTorch native implementation, $180$ for Triton fused, and $250$ for Triton fused (mem save) so as to use approximately $90$ GB of GPU memory for each. The total execution times for Triton fused and Triton fused (mem save) are almost the same, and both are $18$ times faster than the PyTorch native implementation. This drastic acceleration is largely contributed to by accelerating the computation itself through gate fusion and increasing the mini-batch size by three to four times through our memory optimization technique. Therefore, this numerical experiment demonstrates that our theoretical memory analysis is valid, especially for the PyTorch native implementation, and that our proposed method is effective for accelerating the training of QML and VQAs in medium-scale quantum circuits.
@@
Table 2: Results of training the one epoch for minimizing the sum of expectation values at the final state, made by applying a $20$-qubit, $100$-layer ($6,000$ parameters) HEA to $1,000$ randomly initialized states

### 03 Apr 2026 v2

@@ 4.3.2 Numerical experiment for a practical scale QML benchmark
Finally, assuming the training of a practical QML model using 1,0001{,}000 samples of randomly initialized states, we measured the mean training time per epoch and peak memory usage. For the numerical settings, we set $b=10$ for Triton fused and $b=20$ for Triton fused (mem save), based on the previous numerical results with a batch size of one. In addition, we set the mini-batch size to $60$ for Triton fused and $85$ for Triton fused (mem save), to fully use the GPU memory capacity, reaching a peak memory usage of approximately $95$ GB, as shown in Table 1.
As a result, the execution time was 1,2731{,}273 sec for Triton fused and 1,1111{,}111 sec for Triton fused (mem save), indicating that Triton fused (mem save) was approximately $10\%$ faster than Triton fused. Thus, the strategy of storing state vectors in reduced precision accommodates an approximately $1.4$ times larger mini-batch size, thereby enabling more effective utilization of GPU parallel computation resources. This resulted in a throughput improvement, outweighing the overhead of type conversion.
Based on these results, for a $20$-qubit HEA model including 60,00060{,}000 parameters, we estimate the training time to be approximately $15$ to $20$ hours per epoch for a data set containing tens of thousands of training samples, such as MNIST or CIFAR-10. Moreover, since our implementation is compatible with PyTorch’s standard ecosystem, it is easy to use multi-GPU distributed training (e.g., via DistributedDataParallel in PyTorch) to achieve even faster training. Therefore, our proposed method enables training deeper and larger-scale QML models within a realistic timeframe.
@@
Figure 9: (a) Peak memory usage and (b) Execution time, when varying $b$ for a $20$-qubit, 1,0001{,}000-layer (d=1,000d=1{,}000) HEA with a batch size of one
Table 1: Performance comparison for minimizing the sum of expectation values at the final state of an HEA with 2020-qubit, 1,0001{,}000 layers (60,00060{,}000 parameters) for 1,0001{,}000 samples
@@
Triton Fused $60$ $10$ 1,2731{,}273 $94.456$
Triton Fused(mem save) $85$ 2020 1,1111{,}111 $95.378$
@@ 5 Conclusion
Furthermore, by combining our proposed method with gradient checkpointing, we demonstrated the feasibility of training a deep QML model. First, to optimize the checkpoint block size (i.e., the number of HEA layers to be included in the checkpoint block), we measured the mean execution time and peak memory usage with a batch size of one. The results showed that the execution time is almost constant regardless of the checkpoint block size, whereas the peak memory usage varies depending on the checkpoint block size, showing highly consistency with our theoretical analysis. In addition, we found that peak memory usage can be reduced by $30\%$ by temporarily storing the state vectors as bfloat16 converted from float32 (complex64). Using this memory-saving implementation, we trained a $20$-qubit, 1,0001{,}000-layer HEA (60,00060{,}000 parameters) on 1,0001{,}000 samples for one epoch, achieving an execution time of 1,1111{,}111 seconds. Based on this result, we estimate that we could train a $20$-qubit QML model with 60,00060{,}000 parameters on tens of thousands of training samples, such as MNIST or CIFAR-10, within $15$ to $20$ hours per epoch. Also, since our implementation is compatible with standard PyTorch functions, we can further accelerate training by distributing samples over multiple GPUs using utilities in PyTorch, such as DistributedDataParallel. Therefore, our combined approach enables the training of deep QML models on large datasets within a realistic timeframe, which would previously require a supercomputer or a multi-GPU cluster.
@@
Acknowledgements
This work was partially conducted using the Supermicro ARS- [^1] GL-DNHR-LCC (Miyabi-G), equipped with GH200 Grace-Hopper Superchip, at Joint Center for Advanced High Performance Computing (JCAHPC).
@@ Acknowledgements
The code used in this study will be available at GitHub (https://github.com/puyokw/triton_qml/tree/paper).
@@ B.1 The implementation of fused CZ gates and fused CNOT gates
Since a CNOT gate or a CZ gate is a unitary matrix with real values ($U=U^{\dagger}=U^{\top}$) with no variational parameters, we can update the quantum states and adjoint states using the same kernel function for both the forward and backward paths. Here, we describe the implementation for the forward path.
Let $c$ be the index of the control qubit and $t$ be the index of the target qubit. As we explained in Sec. 3.1, applying a CZ gate acting on the $c$-th control qubit and the $t$-th target qubit corresponds to multiplying the probability amplitude $a_{x}$ by 1-1 only when both the $c$-th and $t$-th bits of the index $x$ are $1$. Similarly, for the implementation of fused multiple CZ gates, let $c_{j}$ and $t_{j}$ denote the control and target qubits for the $j$-th gate in the fused operation. Each thread first counts the number of CZ gates where both the $c_{j}$-th and $t_{j}$-th bits of the index $x$ are $1$. If this total count is odd, the thread multiplies the corresponding probability amplitude $a_{x}$ by 1-1.
Specifically, we implemented this using Triton [triton2019], as shown in Listing 1. First, we prepare a bitmask (cz_mask) where only the $c_{j}$-th and $t_{j}$-th bits are set to $1$. For each thread index (offset), we check if the bitwise-AND between the mask and the index (cz_mask & offset) equals the mask itself (cz_mask). By accumulating these boolean evaluation results across all individual CZ gates within the fused operation via an XOR operation to compute the parity, we can determine whether to flip the sign of the amplitude. By applying this calculation to all $2^{n}$ amplitudes, we can update the state vector for the fused CZ gates.
@@
Although we used fused CZ gates in this paper, the efficient implementation of fused CNOT gates is also possible. Here, we describe how to implement a fused CNOT gate with out-of-place operations, since, as we mentioned before, PyTorch’s automatic differentiation requires inputs to be invariant. Also, recall that applying CNOT gate corresponds to swapping the pair of amplitudes between $a_{\ldots 1_{c}\ldots 0_{t}\ldots}$ and $a_{\ldots 1_{c}\ldots 1_{t}\ldots}$ (or between $a_{\ldots 0_{t}\ldots 1_{c}\ldots}$ and $a_{\ldots 1_{t}\ldots 1_{c}\ldots}$), so we can implement it by tracking the index. Specifically, we start from the original index assigned to each thread and trace it to determine the final destination index. As described in Listing 2, for the $j$-th CNOT gate in the fused operation, if the $c_{j}$-th bit of the current index is $1$, we update the current index (out_idx) by applying a bitwise XOR operation between the current index and a mask (target_mask) that has only the $t_{j}$-th bit set to $1$. By repeating this procedure across all individual CNOT gates within the fused operation, we obtain the final destination index. By writing the original probability amplitude to the resulting destination index (out_idx) in a newly allocated vector, we can update the quantum states as an out-of-place operation.
@@ Appendix C Memory behavior analysis of the PyTorch native implementation using $100$-layer HEA
In this section, to complement the experiments in Sec. 4.3, we describe the results of a classical simulation for $100$ HEA layers to verify the memory behavior of the PyTorch native implementation, which could not be evaluated in the large-scale numerical experiments in Sec. 4.3 due to excessive memory requirements. Specifically, we verify our analytical model described in Sec. 3.4, especially for the PyTorch native implementation, using a smaller model consisting of $100$ HEA layers with $20$ qubits and 6,0006{,}000 parameters. Here again, we performed the task of minimizing the sum of expectation values for the observable $O=IXYZ\ldots IXYZ$ (consisting of $5$ repetitions of $IXYZ$) at the final states, generated by applying the quantum circuit to random initial quantum states for a $20$-qubit system.
@@
which is obtained by substituting $l_{\text{var}}=60,\alpha=9,d=100,M_{\text{sv}}=8\times 2^{20}$ into Eq. (7), noting that $\lceil\frac{l_{\text{var}}}{\alpha}\rceil=7$. Based on this equation, the theoretical optimal block size is $b^{*}=3.7$, which closely matches our empirical observation where the minimum peak memory usage ($528$ MB) was recorded at $b=4$. For the Triton fused (mem save) implementation, the theoretical peak memory usage is estimated by
@@
So, the theoretical optimal block size for the Triton fused (mem save) implementation is $b^{*}=5.3$, which again aligns well with the measured minimum peak memory usage ($406$ MB) observed at $b=5$.
Overall, as shown in Fig. 11(a), we can see that the theoretical values for peak memory usage are highly consistent with the measured values across all implementations. Note that a slight discrepancy between theoretical and measured values is observed when memory usage is below 1 GB, primarily due to the memory overheads such as CUDA contexts.
@@
Comparing the training time for $1$ epoch: Here, we compare the training times among the three implementations. For the checkpoint setting, based on the results of the above numerical experiments, we set the checkpoint block size $b$ to $b=1$ (every HEA layer) for the PyTorch native implementation, $b=4$ for Triton fused, and $b=5$ HEA layers for Triton fused (mem save). We performed the task of minimizing the sum of expectation values for the observable $O=(IXYZ)^{\otimes 5}$ at the final state, generated by applying $100$ HEA layers to 1,0001{,}000 randomly initialized quantum states with a $20$-qubit system, over $1$ epoch.
As shown in Table 2, we set the mini-batch size to $60$ for the PyTorch native implementation, $180$ for Triton fused, and $250$ for Triton fused (mem save) so as to use approximately $90$ GB of GPU memory for each. The total execution times for Triton fused and Triton fused (mem save) are almost the same, and both are $18$ times faster than the PyTorch native implementation. This drastic acceleration is largely achieved by reducing memory access overhead and increasing the mini-batch size by three to four times through gate fusion and our memory-saving technique. Therefore, this numerical experiment demonstrates that our theoretical memory analysis is valid, especially for the PyTorch native implementation, and that our proposed method is effective for accelerating the training of QML and VQAs in medium-scale quantum circuits.
@@
Table 2: Results of training a $20$-qubit, $100$-layer (6,0006{,}000 parameters) HEA for one epoch to minimize the sum of expectation values over 1,0001{,}000 randomly initialized states